Exercise 4.3 Given P130.0 and P135.2. calculata the inflation rate in period t s Center your es an inleer
Answer #1

Inflation rate in period t = [(Pt -Pt-1) / Pt-1 ]100 = [(135.2-130)/ 130] 100

= (5.2/130)100 = (0.04)(100)= 4%